OpenStudy (anonymous):

The results of a class election are shown in the table below. There were a total of 600 votes. For each candidate, write a fraction that shows the number of votes and the total number of votes. Then write each fraction as a decimal. jeremy- 152 bina-104 jun-136 gregory-208

OpenStudy (anonymous):

jeremy = 152/600 = 0.2533 bina = 104/600 = 0.1733 jun = 136/6000.2266 gregory = 208/600 =0.3466